# Reformulating Domain Adaptation of Large Language Models as Adapt-Retrieve-Revise

### v1

@@ 4.1.1 Retrieving A Query or Retrieving an Answer?
We further compare the query-based and answer-based retrieval on a public Similar Case Retrieval task. This task aims to select similar legal judgments based on the query from the candidates with a query the case brief given. As shown in Table 2, we compare two setups: 1) using the original query to retrieve, 2) using the legal 7B LLM to complete a whole legal judgment document given the brief query, and then retrieving. We follow the original task repository for the other settings: https://github.com/myx666/LeCaRD. The results show that on each metric, the answer-based retrieval works better, indicating that using the generated answer provides a more robust retrieval.
@@
Setup Precision@$5$ Precision@$10$ MAP
Query-based 42.08 41.95 47.8
Answer-based 45.19 (+3.11) 42.08 (+0.13) 49.49 (+1.69)
@@ 4.2 Case analysis of the improvements after the GPT-4 revision
  • Revise hallucinations in the evidence: even the domain-adapted LLM can provide evidence from its learned legal knowledge; the hallucination remains to some degree, such as the wrong clause index, even the law name and rationale are roughly correct. Since the rationale content is accurate, the answer-based retrieval will search for the correct evidence, and the revision by GPT-4 will solve the hallucination to produce a more robust response.
  • Choose correct evidence: In a more significant scenario, even though the 7B legal model’s answers might contain partial hallucinatory content, the retrieval component can still possibly identify correct evidence through the partially correct descriptions in the rationale generation. During the revision stage, GPT-4 could assess the correct evidence, leading to the generation of correct answers.
Figure 6: Case studies on the GPT-4 revision.
@@ 5 Related Work
5.1 Tasks in the Chinese Legal Domain
The rapid advancements in LLMs have significantly impacted various domains, including the legal industry. This gives rise to the occurrence of legal datasets, such as the Challenge of AI in Law (CAIL, http://cail.cipsc.org.cn/index.html), LeCaRD (Ma et al., 2021), JEC-QA (Zhong et al., 2020) and EQUALS (Chen et al., 2023). These datasets cover document classification, summarization, question answering, information extraction, similar document retrieval, and other popular NLP tasks in the Chinese legal domain. To the best of our knowledge, this paper is the first work to exam the zero-shot performances on these legal datasets.
@@ 5.3 Retrieval-augmented inference
In scenarios where language models are confronted with tasks necessitating an infusion of external knowledge, a retriever mechanism can be used to provide evidence. The Retrieval-Augmented Generation (RAG) (Lewis et al., 2020b) system incorporates a BERT-based (Devlin et al., 2019) Document Retrieval Process (DRP) and utilizes BART (Lewis et al., 2020a) for answer generation. Analogously, the EMDR2 (Yu et al., 2023) employs the expectation-maximization algorithm to account for multiple retrieved documents. The Atlas (Izacard et al., 2022) builds upon the EMDR2 framework, and by synergistically training the retriever and reader components, it demonstrates few-shot learning capabilities commensurate with the 540B PalM (Chowdhery et al., 2022). RETRO (Borgeaud et al., 2022) benefits from retrieval mechanisms on expansive corpora during its pre-training phase and exhibits performance in close alignment with those of GPT-3 (Brown et al., 2020b).
6 Conclusions and Future Discussions
In this paper, we reformulate the zero-shot domain content generation of large language models as an adapt-retrieve-revise procedure. This approach combines the merits of efficiently performing continual training of a smaller 7B LLM for domain adaptation, robustly retrieving the supporting evidence from an external knowledge base, and effectively leveraging the evidence-assessing and revision capabilities of GPT-4. Our method significantly enhances the zero-shot performance of GPT-4 on four Chinese legal tasks.

### v3

@@ 5.1.1 Retrieving A Query or Retrieving an Answer?
We further compare the query-based and answer-based retrieval on a public Similar Case Retrieval task. This task aims to select similar legal judgments based on the query from the candidates with a query the case brief given. As shown in Table 3, we compare two setups: 1) using the original query to retrieve, 2) using the legal 7B LLM to complete a whole legal judgment document given the brief query, and then retrieving. We follow the original task repository for the other settings [^3] https://github.com/myx666/LeCaRD. The results show that on each metric, the answer-based retrieval works better, indicating that using the generated answer provides a more robust retrieval.
Setup Precision@$5$ Precision@$10$ MAP
Query-based 42.1 42.0 47.8
Answer-based 45.2 (+3.1) 42.1 (+0.1) 49.5 (+1.7)
@@
Figure 6: Case studies on the GPT-4 revision. The examples are translated from Chinese to English for the demonstration purpose.
@@ 5.2 Case Analysis of the Improvements after the GPT-4 Revision
  • Revising hallucinations in the evidence: even the domain-adapted LLM can provide evidence from its learned legal knowledge; the hallucination remains to some degree, such as the wrong clause index, even the law name and rationale are roughly correct. Since the rationale content is accurate, the answer-based retrieval will search for the correct evidence, and the revision by GPT-4 will solve the hallucination to produce a more robust response.
  • Choosing correct evidence: In a more significant scenario, even though the 7B legal model’s answers might contain partial hallucinatory content, the retrieval component can still possibly identify correct evidence through the partially correct descriptions in the rationale generation. During the revision stage, GPT-4 could assess the correct evidence, leading to the generation of correct answers.
Figure 7: Comparison of iterations on LegalQA dataset.
@@ 6 Related Work
6.1 Chinese Legal Domain Tasks
The rapid advancements in LLMs have significantly impacted various domains, including the legal industry. This gives rise to the occurrence of legal datasets, such as the Challenge of AI in Law (CAILf)444 http://cail.cipsc.org.cn/index.html, LeCaRD (Ma et al., 2021), JEC-QA (Zhong et al., 2020) and EQUALS (Chen et al., 2023). These datasets cover document classification, summarization, question answering, information extraction, similar document retrieval, and other popular NLP tasks in the Chinese legal domain. To the best of our knowledge, this paper is the first work to exam the zero-shot performances on these legal datasets.
@@ 6.3 Retrieval-augmented Inference
In scenarios where language models are confronted with tasks necessitating an infusion of external knowledge, a retriever mechanism can be used to provide evidence. The Retrieval-Augmented Generation (RAG) (Lewis et al., 2020b) system incorporates a BERT-based (Devlin et al., 2019) Document Retrieval Process (DRP) and utilizes BART (Lewis et al., 2020a) for answer generation. Analogously, the EMDR2 (Yu et al., 2023) employs the expectation-maximization algorithm to account for multiple retrieved documents. The Atlas (Izacard et al., 2022b) builds upon the EMDR2 framework, and by synergistically training the retriever and reader components, it demonstrates few-shot learning capabilities commensurate with the 540B PalM (Chowdhery et al., 2022). RETRO (Borgeaud et al., 2022) benefits from retrieval mechanisms on expansive corpora during its pre-training phase and exhibits performance in close alignment with those of GPT-3 (Brown et al., 2020b).
7 Conclusions
In this paper, we reformulate the zero-shot domain content generation of large language models as an adapt-retrieve-revise procedure. This approach combines the merits of efficient continual training of a smaller 7B LLM for domain adaptation, robustly retrieving the supporting evidence from an external knowledge base, and effectively leveraging the evidence-assessing and revision capabilities of GPT-4. Our proposal substantially enhances the zero-shot performance on the Chinese legal tasks.
8 Limitations
@@ A.1 Legal Instruction Tuning
A.2 Retrieval Modules
we leveraged multilingual E5-large, which is the SOTA family of text embeddings, which has been reported to outperform BM25, Contriever (Izacard et al., 2022a), and GPT embeddings (OpenAI, 2023b). Since the improvements of our method are consistent and substantial (+33.3% vs vanilla, +15.4%/23.9% vs retrieval baselines), we believe these gaps have shown sufficient robustness in our proposal.
However, we agree that including more established retrieval modules can enhance the robustness of our findings. Therefore, we added extra experiments and compared them with the current SOTA Chinese retrieval module CoROM following (Qiu et al., 2022) on the LegalQA dataset, the same setting as in Section 5.1. Table 4 shows the comparison results. We find that: (1) multilingual e5-large has a competitive performance with CoROM on query-based retrieval and vastly outperforms CoROM on answer-based retrieval; (2) for both modules, the answer-based retrieval primarily improves the retrieval quality than the query-based setting.
Retriever Retrieval Top-1 Top-5 Top-10
Multilingual E5-large Query-based 45.8 61.3 70.5
Multilingual E5-large Answer-based 65.3 84.5 88.5
CoROM Query-based 47.5 60.8 71.5
CoROM Answer-based 58.8 72.5 80.5
Table 4: Results of different retrieval seteps.